# Monte Carlo Neural PDE Solver for Learning PDEs via Probabilistic Representation

### v4

@@ 1 Introduction
Beyond these achievements, the function-to-function neural PDE solvers have risen as another new paradigm for simulating physical systems, which leverage neural networks as surrogate models to approximate the solutions of a family of PDEs [74, 86, 46, 42]. Along this direction, several studies have proposed diverse network architectures for neural PDE solvers [74, 41, 46, 7]. These solvers can be trained using supervised [74, 46, 41] or unsupervised approaches [86, 89, 43], employing pre-generated data or PDE information to construct training targets, respectively. The unsupervised training approach is essential for DL-based PDE solvers, particularly in scenarios with limited available or high-quality data. To address this issue, some studies [86, 78, 43] borrow techniques from classical numerical solvers to construct training targets. For instance, the MAC grid physics-constrained network and its 3D extension [86, 87], low-rank decomposition network (LordNet) [78] and physics-informed neural operator (PINO) [43] integrate finite difference or pseudo-spectral methods with neural networks during the training stage. However, these traditional Eulerian methods require fine meshes or step sizes for stable simulations. Therefore, the performance and efficiency of corresponding neural PDE solvers are also limited by time and space discretization, particularly when handling high spatiotemporal variations. Furthermore, these methods often necessitate additional loss functions to enforce boundary conditions, introducing extra complexity in hyperparameter tuning and computational demands.
@@
3. Beyond comparisons with unsupervised learning methods, we comprehensively compare MCNP Solver with other widely-used PDE solvers, including the classical Eulerian solver, Monte Carlo methods, and the supervised training approach. We delve into a detailed discussion on each method’s strengths, weaknesses, and application scopes.
@@ 3.1 Preliminary
where $\boldsymbol{x}\in\Omega\subset\mathbb{R}^{d}$ and $t$ denote the $d$-dimensional spatial variable and the time variable, $\boldsymbol{\beta}[u](\boldsymbol{x},t)\in\mathbb{R}^{d}$ is a vector-valued mapping from $(u,\boldsymbol{x},t)$ to $\mathbb{R}^{d}$, $\kappa\in\mathbb{R}^{+}$ is the diffusion parameter, and $f(\boldsymbol{x},t)\in\mathbb{R}$ denotes the force term. Many well-known PDEs, such as the Allen–Cahn and Navier-Stokes equations, can be viewed as a special form of Eq. 1.
@@ 3.3.1 The Simulation of Convection Process
$\boldsymbol{\xi}_{p,t}^{c}$ $=\boldsymbol{\xi}_{p,t+\Delta t}+\frac{1}{2}\left(\boldsymbol{\beta}_{p,t}^{c,1}+\boldsymbol{\beta}_{p,t}^{c,2}\right),$
where $u$ is the output of the MCNP Solver in each iteration.
@@ 4 Experiments
Table 1: 1D convection-diffusion equation with varying $\beta$ and $\kappa$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined.
@@ 4.1.3 Results
Fig. 4 illustrates the predicted $u$ for MCNP-10 from $t=0.2$ to $t=2.0$ for E1-E6, respectively. When comparing the cases with $\beta=0.01$ and $0.1$, the effect of material advection becomes more remarkable with larger $\beta$, with peaks and troughs of the heat flows more noticeably shifting towards the negative $x$-axis direction. Keeping $\beta$ as a constant, as $\kappa$ increases, we observe that the diffusion effect gradually governs the motion of the material, causing the peaks and troughs to decay rapidly. In particular, E4 exhibits the most significant convection effect due to the Péclet number ($\operatorname{Pe}\propto\beta/\kappa$) reaching maximum in this example. When $\operatorname{Pe}$ number is high, the convection effect is dominant, and the diffusion effect can be disregarded, which means that the macroscopic motion of the heat flow has a much greater impact than the random motion of particles. Table 1 presents each method’s performance and computational cost in the 1D convection-diffusion equation. Among all unsupervised neural PDE solvers, including PI-DeepONet-(M) and PINO-10/20/100, the MCNP-10 performs the best on most tasks, particularly for cases with large convection or diffusion rates. For PINO, its accuracy is not robust with step size, indicating that the pseudo-spectral method cannot provide sufficiently accurate training targets with a coarser partition of the time interval. Additionally, the accuracy of PINO in E4 declines most rapidly with increasing temporal step size. Such phenomenons happen because the $\operatorname{Pe}$ number in this experiment is significantly higher than in other experiments; hence, the conventional Eulerian method requires a finer time step to maintain simulation accuracy. For PINO-100, despite achieving comparable results on E1 and E5, its training cost is eight times more than that of MCNP-10. It is also interesting to note that MCNP-10 performs better than MCNP-20 in Table 1. The underlying reason is that the drift coefficient $\beta$ and the forcing $f$ are constants in Eq. 19, and thus, the discrete errors can be eliminated when simulating the trajectories of particles for Feynman-Kac law. As a result, the setting $\Delta t=2/10$ is already sufficient to provide accurate training signals for the MCNP Solver, while further refining the step size would only add additional fitting and generalization burdens.
@@
$\frac{\partial u(x,t)}{\partial t}=$ $0.01\Delta u(x,t)+u(x,t)-u(x,t)^{3},$ (20)
@@ 4.2.1 Experimental Settings
In this section, we select two different $N$ in $\{5,10\}$ and two kinds of boundary conditions, to evaluate the performance of different methods in handling spatial variations and varying boundary conditions. These four experimental settings are denoted E1-E4 with $(N,\operatorname{Boundary})=(5,\operatorname{Dirichlet}),(10,\operatorname{Dirichlet}),(5,\operatorname{Neumann}),(10,\operatorname{Neumann})$, respectively. We divide the spatial domain $[0,1]$ into $65$ uniform grid elements for all experiments.
@@ 4.2.3 Results
Figure 5: Simulation of 1D Allen-Cahn equation. The prediction result (Left) and point-wise error (Right) of MCNP-100 for an example in E1-E4. The x-axis and y-axis represent spatial coordinates and the predicted values (point-wise error).
Fig. 5 illustrates the predicted $u$ for MCNP-100 from $t=0.1$ to $t=1.0$ for E1 through E4, respectively. The simulation of the Allen-Cahn equation is more challenging than the convection-diffusion equation due to its nonlinear nature. We can observe that higher $N$ values can introduce more obvious spatial variations, particularly near the initial time. In the experiments with Dirichlet boundary conditions, the system values decay rapidly near the boundary, whereas in the case of Neumann conditions, they maintain a relatively stable variation. From a microscopic perspective, this occurs because the particles are absorbed after hitting the Dirichlet boundary, while they are reflected after colliding with the Neumann boundary. Unlike other baseline methods, MCNP Solver does not rely on additional loss functions to encode the boundary conditions. Instead, the MCNP Solver can naturally satisfy these boundary conditions by choosing the corresponding transition probability during its random walk. The simulation results in Fig. 5 demonstrate that the MCNP Solver effectively meets the Dirichlet and Neumann boundary conditions when solving the Allen-Cahn equations. Table 2 presents each method’s performance and computational cost on the 1D Allen-Cahn equation. The results of PI-DeepONet indicate that the PINN loss cannot efficiently handle high-frequency components, which has also been observed in previous literature [35, 91]. Although PI-DeepONet-M mitigates this issue via adaptive sampling, there is still a gap in precision when solving high-frequency problems compared to other methods. Among all unsupervised methods, only MCNP-100 achieves a relative error lower than 1% on all experiments and metrics. Moreover, the performances of MCNP-20 are generally comparable with PINO-100 while taking only around 21% of the training time, demonstrating the advantages of neural Monte Carlo loss for learning PDEs in an unsupervised manner.
Table 2: 1D Allen-Cahn equation with varying $N$ and boundary conditions. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined.
@@ 4.3.1 Experimental Settings
Table 3: 2D Navier-Stokes equation with varying $\nu$ and forcing. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined. The training time for MCNP varies among the six experiments due to the interpolation trick.
@@ 4.3.3 Results
Fig. 6 shows the predicted vorticity field $\omega$ of a learned MCNP Solver from $t=2$ to $t=10$ in E1-E6, respectively. Compared to the case with $\nu=10^{-4}$, we observe that the vorticity field has more remarkable peak and trough values with more intricate details when $\nu=10^{-5}$, which implies that the fluid elements are rotating at higher speeds. This instability is caused by the nonlinear convection terms gradually taking control of the motion of fluids. Additionally, the external forcing $f$ significantly influences the trend of fluid motion. When the external forcing is zero, the fluid motion primarily arises from internal dynamical processes within the system, such as the interaction between viscous and inertial forces, resulting in a relatively slow change in the vorticity field over time. When an external forcing drives the system, the corresponding vorticity field changes markedly in response to the forcing, leading to a faster evolution of the vorticity field. In particular, when driven by a Kolmogorov forcing, numerous new vortices of different scales are generated during the flow, making the fluid simulation more challenging. More simulation results of other baseline methods can be seen in the Appendix  IV. Table 3 presents each method’s performance and computational cost on the 2D Navier-Stokes equations. Compared to other unsupervised training methods, the MCNP Solver achieved the lowest relative error for all tasks and metrics. Unlike PINO, whose performance is highly affected by step size, especially for cases with $\nu=10^{-5}$, the MCNP Solver can maintain stability with respect to step size. It is worth mentioning that MCNP-10 outperforms PINO-100 on all metrics while only taking 13.1-14.5% of the training time. Such an advantage stems from the fact that the Lagrangian method can be better adapted to coarse step size, as discussed in [28, 34].
@@
As shown in Eq. 22, calculating the fractional Laplacian operator is challenging due to its singularity and non-local properties. However, we can efficiently simulate it from the probability perspective, where we only need to replace the Brownian motion in Eq. 4 with the $\alpha$-stable Lévy process [33, 98, 97]. The initial states $u(\boldsymbol{x},0)$ are generated from the functional space $\mathcal{H}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}\sin(\pi n\|\boldsymbol{x}\|_{2}^{2}):a_{n}\sim\mathbb{U}(0,1)\}$, and $N$ represents the maximum frequency of the functional space.
@@ 4.4.1 Experimental Settings
We select four different $\alpha$ in $\{0.5,1.0,1.5,2.0\}$ to evaluate the performance of MCNP Solver in handling varying fractional coefficients. These four experiments are denoted as E1-E4. We divide the time interval $[0,1]$ into $100$ uniform intervals. We set the maximum frequency $N$ as 10. The test data is generated from the fine-grained Monte Carlo methods with $\Delta t=10^{-3}$, which samples $10^{4}$ particles when simulating the SDEs. Because Eq. 22 is defined on a unit disk, we utilize the DeepONet as a backbone model and train MCNP Solver in a mesh-free regime.
@@
Table 4: 2D fractional diffusion equation with varying $\alpha$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs.
@@
E1 ($\alpha=0.5$) 3.812 $\pm$0.219 4.322 $\pm$0.253 1.260 0.325
E2 ($\alpha=1.0$) 4.070 $\pm$0.190 5.057 $\pm$0.336 1.260 0.325
E3 ($\alpha=1.5$) 2.885 $\pm$0.587 3.781 $\pm$0.734 1.260 0.325
E4 ($\alpha=2.0$) 3.808 $\pm$1.263 5.021 $\pm$1.332 1.260 0.325

### v5

@@ Abstract
Index Terms:
Neural PDE solver, Monte Carlo method, Feynman-Kac formula, AI for PDE.
@@ 1 Introduction
Beyond these achievements, the function-to-function neural PDE solvers have risen as another new paradigm for simulating physical systems, which leverage neural networks as surrogate models to approximate the solutions of a family of PDEs [76, 87, 48, 43]. Along this direction, several studies have proposed diverse network architectures for neural PDE solvers [76, 42, 48, 7, 98]. These solvers can be trained using supervised [76, 48, 42] or unsupervised approaches [87, 90, 45], employing pre-generated data or PDE information to construct training targets, respectively. The unsupervised training approach is essential for DL-based PDE solvers, particularly in scenarios with limited available or high-quality data. To address this issue, some studies [87, 80, 45] borrow techniques from classical numerical solvers to construct training targets. For instance, the MAC grid physics-constrained network and its 3D extension [87, 88], low-rank decomposition network (LordNet) [80] and physics-informed neural operator (PINO) [45] integrate finite difference or pseudo-spectral methods with neural networks during the training stage. However, these traditional Eulerian methods require fine meshes or step sizes for stable simulations. Therefore, the performance and efficiency of corresponding neural PDE solvers are also limited by time and space discretization, particularly when handling high spatiotemporal variations. Furthermore, these methods often necessitate additional loss functions to enforce boundary conditions, introducing extra complexity in hyperparameter tuning and computational demands.
@@
3. Beyond comparisons with unsupervised learning methods, we comprehensively compare MCNP Solver with other widely-used PDE solvers, including the Eulerian solver, Monte Carlo methods, and the supervised training approach. We delve into a detailed discussion on each method’s strengths, weaknesses, and application scopes.
@@ 3.1 Preliminary
where $\boldsymbol{x}\in\Omega\subset\mathbb{R}^{d}$ and $t$ respectively denote the spatial variable and the temporal variable, $\boldsymbol{\beta}[u](\boldsymbol{x},t)\in\mathbb{R}^{d}$ is a vector-valued mapping from $(u,\boldsymbol{x},t)$ to $\mathbb{R}^{d}$, $\kappa\in\mathbb{R}^{+}$ is the diffusion parameter, and $f(\boldsymbol{x},t)\in\mathbb{R}$ denotes the force term. $\nabla u$ and $\Delta u$ demote the gradient and Laplacian of $u$, respectively. $u_{0}(\boldsymbol{x})$ represents the initial condition. Many well-known PDEs, such as the Allen–Cahn and Navier-Stokes equations, can be viewed as a special form of Eq. 1.
@@ 3.3.1 The Simulation of Convection Process
$\boldsymbol{\xi}_{p,t}^{c}$ $=\boldsymbol{\xi}_{p,t+\Delta t}+\frac{1}{2}\left(\boldsymbol{\beta}_{p,t}^{c,1}+\boldsymbol{\beta}_{p,t}^{c,2}\right).$
For the cases in which the drift term $\boldsymbol{\beta}$ depends on solution $u$, we utilize the output of MCNP Solver to approximate $\boldsymbol{\beta}[u]$ accordingly.
@@ 4 Experiments
TABLE I: 1D convection-diffusion equation with varying $\beta$ and $\kappa$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined.
@@ 4.1.3 Results
Fig. 4 illustrates the predicted $u$ for MCNP-10 from $t=0.2$ to $t=2.0$ for E1-E6, respectively. When comparing the cases with $\beta=0.01$ and $0.1$, the effect of material advection becomes more remarkable with larger $\beta$, with peaks and troughs of the heat flows more noticeably shifting towards the negative $x$-axis direction. Keeping $\beta$ as a constant, as $\kappa$ increases, we observe that the diffusion effect gradually governs the motion of the material, causing the peaks and troughs to decay rapidly. In particular, E4 exhibits the most significant convection effect due to the Péclet number ($\operatorname{Pe}\propto\beta/\kappa$) reaching maximum in this example. When $\operatorname{Pe}$ number is high, the convection effect is dominant, and the diffusion effect can be disregarded, which means that the macroscopic motion of the heat flow has a much greater impact than the random motion of particles. Table I presents each method’s performance and computational cost in the 1D convection-diffusion equation. Among all unsupervised neural PDE solvers, including PI-DeepONet-(M) and PINO-10/20/100, the MCNP-10 performs the best on most tasks, particularly for cases with large convection or diffusion rates. For PINO, its accuracy is not robust with step size, indicating that the pseudo-spectral method cannot provide sufficiently accurate training targets with a coarser partition of the time interval. Additionally, the accuracy of PINO in E4 declines most rapidly with increasing temporal step size. Such phenomenons happen because the $\operatorname{Pe}$ number in this experiment is significantly higher than in other experiments; hence, the conventional Eulerian method requires a finer time step to maintain simulation accuracy. For PINO-100, despite achieving comparable results on E1 and E5, its training cost is eight times more than that of MCNP-10. It is also interesting to note that MCNP-10 performs better than MCNP-20 in Table I. The underlying reason is that the drift coefficient $\beta$ and the forcing $f$ are constants in Eq. 19, and thus, the discrete errors can be eliminated when simulating the trajectories of particles for Feynman-Kac law. As a result, the setting $\Delta t=2/10$ is already sufficient to provide accurate training signals for the MCNP Solver, while further refining the step size would only add additional fitting and generalization burdens.
@@
$\frac{\partial u(x,t)}{\partial t}=$ $\epsilon\Delta u(x,t)+u(x,t)-u(x,t)^{3},$ (20)
@@ 4.2.1 Experimental Settings
Firstly, we fix $\epsilon$ as 0.01, and select two different $N$ in $\{5,10\}$ and two kinds of boundary conditions to evaluate the performance of different methods in handling spatial variations and varying boundary conditions. Secondly, when $\epsilon$ tends to zero, the interface layers become thinner and sharper, leading to long-lived metastable structures [15]. To observe such phenomena, we choose $\epsilon$ as 0.0001 with $N$ in $\{5,10\}$, respectively. Therefore, these six experimental settings are denoted E1-E6 with $(\epsilon,N,\operatorname{Boundary})=(0.01,5,\operatorname{D}),(0.01,10,\operatorname{D}),(0.01,5,\operatorname{N}),\\ (0.01,10,\operatorname{N}),(0.0001,5,\operatorname{D}),(0.0001,10,\operatorname{D})$111”D” represents the Dirichlet boundary condition and ”N” represents the Neumann boundary condition., respectively. We divide the spatial domain $[0,1]$ into $65$ uniform grid elements for all experiments.
@@ 4.2.3 Results
Figure 5: Simulation of 1D Allen-Cahn equation. The prediction result (Left) and point-wise error (Right) of MCNP-100 for an example in E1-E6. The x-axis and y-axis represent spatial coordinates and the predicted values (point-wise error).
Fig. 5 illustrates the predicted $u$ for MCNP-100 from $t=0.1$ to $t=1.0$ for E1 through E6, respectively. The simulation of the Allen-Cahn equation is more challenging than the convection-diffusion equation due to its nonlinearity. We can observe that higher $N$ values can result in more obvious spatial variations, particularly near the initial time. In the experiments with Dirichlet boundary conditions, the system values decay rapidly near the boundary, whereas in the case of Neumann conditions, they maintain a relatively stable variation. From a microscopic perspective, this occurs because the particles are absorbed after hitting the Dirichlet boundary, while they are reflected after colliding with the Neumann boundary. Unlike other baseline methods, MCNP Solver does not rely on additional loss functions to encode the boundary conditions. Instead, the MCNP Solver can naturally satisfy these boundary conditions by choosing the corresponding transition probability during its random walk. The simulation results in Fig. 5 demonstrate that the MCNP Solver effectively meets the Dirichlet and Neumann boundary conditions when solving the Allen-Cahn equations. It can be seen in Fig. 5 that when $\epsilon=0.0001$, the phase field $u(x)$ oscillates rapidly between positive and negative values, highlighting the sharpness of the interface layers due to the very small $\epsilon$. These oscillations are caused by the equation striving to minimize the energy, leading to high-frequency changes in the solution near the interface to accommodate the small $\epsilon$ [15]. Table II presents each method’s performance and computational cost on the 1D Allen-Cahn equation. The results of PI-DeepONet indicate that the PINN loss cannot efficiently handle high-frequency components, which has also been observed in previous literature [36, 92]. Although PI-DeepONet-M mitigates this issue via adaptive sampling, there is still a gap in precision when solving high-frequency problems compared to other methods. Among all unsupervised methods, only MCNP-100 achieves a relative error lower than 1% on all experiments and metrics. Moreover, the performance of MCNP-20 is generally comparable to that of PINO-100 while taking only around 21% of the training time, demonstrating the advantages of neural Monte Carlo loss to learn PDEs unsupervised.
TABLE II: 1D Allen-Cahn equation with varying $\epsilon$, $N$ and boundary conditions. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined.
@@
E5 $\epsilon=0.0001$ $N=5$ Dirichlet PINO-20 0.735 $\pm$0.162 1.445 $\pm$0.275 0.135 0.140
PINO-100 0.330 $\pm$0.004 0.842 $\pm$0.003 0.627 0.140
PI-DeepONet 8.251 $\pm$2.665 15.853 $\pm$4.519 0.431 0.219
PI-DeepONet-M 3.781$\pm$ 1.943 4.724 $\pm$2.568 1.004 0.219
MCNP-20 0.285 $\pm$0.010 0.593 $\pm$0.014 0.156 0.140
MCNP-100 0.303 $\pm$0.018 0.594 $\pm$0.047 0.651 0.140
E6 $\epsilon=0.0001$ $N=10$ Dirichlet PINO-20 3.605 $\pm$1.030 7.181 $\pm$1.636 0.135 0.140
PINO-100 1.817 $\pm$0.008 4.114 $\pm$0.042 0.627 0.140
PI-DeepONet 15.009 $\pm$2.398 25.870 $\pm$3.776 0.431 0.219
PI-DeepONet-M 9.542$\pm$ 1.058 12.357 $\pm$1.569 1.004 0.219
MCNP-20 1.459 $\pm$0.007 2.957 $\pm$0.011 0.156 0.140
MCNP-100 1.487 $\pm$0.090 3.083 $\pm$0.230 0.651 0.140
@@ 4.3.1 Experimental Settings
TABLE III: 2D Navier-Stokes equation with varying $\nu$ and forcing. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs for baseline methods and MCNP Solver. The best results are marked in bold, and the second best results are underlined. The training time for MCNP varies among the six experiments due to the interpolation trick.
@@ 4.3.3 Results
Fig. 6 shows the predicted vorticity field $\omega$ of a learned MCNP Solver from $t=2$ to $t=10$ in E1-E6, respectively. Compared to the case with $\nu=10^{-4}$, we observe that the vorticity field has more remarkable peak and trough values with more intricate details when $\nu=10^{-5}$, which implies that the fluid elements are rotating at higher speeds. This instability is caused by the nonlinear convection terms gradually taking control of the motion of fluids. Additionally, the external forcing $f$ significantly influences the trend of fluid motion. When the external forcing is zero, the fluid motion primarily arises from internal dynamical processes within the system, such as the interaction between viscous and inertial forces, resulting in a relatively slow change in the vorticity field over time. When an external forcing drives the system, the corresponding vorticity field changes markedly in response to the forcing, leading to a faster evolution of the vorticity field. In particular, when driven by a Kolmogorov forcing, numerous new vortices of different scales are generated during the flow, making the fluid simulation more challenging. More simulation results of other baseline methods can be seen in the Appendix  V. Table III presents each method’s performance and computational cost on the 2D Navier-Stokes equations. Compared to other unsupervised training methods, the MCNP Solver achieved the lowest relative error for all tasks and metrics. Unlike PINO, whose performance is highly affected by step size, especially for cases with $\nu=10^{-5}$, the MCNP Solver can maintain stability with respect to step size. It is worth mentioning that MCNP-10 outperforms PINO-100 on all metrics while only taking 13.1-14.5% of the training time. Such an advantage stems from the fact that the Lagrangian method can be better adapted to coarse step size, as discussed in [29, 35].
@@
As shown in Eq. 22, calculating the fractional Laplacian operator is challenging due to its singularity and non-local properties. However, we can efficiently simulate it from the probability perspective, where we only need to replace the Brownian motion in Eq. 4 with the $\alpha$-stable Lévy process [34, 99, 100]. The initial states $u(\boldsymbol{x},0)$ are generated from the functional space $\mathcal{H}_{N}\triangleq\{\sum_{n=1}^{N}a_{n}(1-\|\boldsymbol{x}\|_{2}^{2})^{n+1}:a_{n}\sim\mathbb{U}(0,1)\}$, and $N$ represents the maximum degree of the functional space.
@@ 4.4.1 Experimental Settings
We select four different $\alpha$ in $\{0.5,1.0,1.5,2.0\}$ to evaluate the performance of MCNP Solver in handling varying fractional coefficients. These four experiments are denoted as E1-E4. We divide the time interval $[0,1]$ into $100$ uniform intervals. We set the maximum degree $N$ as 10. Because Eq. 22 is defined on a disk, we use DeepONets as backbone models and train MCNP Solver in a mesh-free regime.
@@
TABLE IV: 2D fractional diffusion equation with varying $\alpha$. Relative $\ell_{2}$ errors ($E_{\ell_{2}}$), relative $\ell_{\infty}$ errors ($E_{\ell_{\infty}}$) and computational costs.
@@
E1 ($\alpha=0.5$) 1.139 $\pm$0.145 1.689 $\pm$0.320 0.139 0.325
E2 ($\alpha=1.0$) 1.538 $\pm$0.616 1.846 $\pm$1.015 0.139 0.325
E3 ($\alpha=1.5$) 1.424 $\pm$0.266 1.386 $\pm$0.315 0.139 0.325
E4 ($\alpha=2.0$) 0.977 $\pm$0.045 1.300 $\pm$0.132 0.139 0.325